OpenStudy (bossimbacon):

Correctly simplify the expression 8 ÷ -2/3? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

think of 8 as 8/1 and use the rule I posted in your last question

OpenStudy (bossimbacon):

would it be -4/3 then

jimthompson5910 (jim_thompson5910):

\[\Large 8 \div -\frac{2}{3}\] \[\Large \frac{8}{1} \div -\frac{2}{3}\] \[\Large \frac{8}{1} \times -\frac{3}{2}\] I'll let you finish up

OpenStudy (bossimbacon):

12? @jim_thompson5910 all simplifyed

jimthompson5910 (jim_thompson5910):

-12, you forgot the negative sign

OpenStudy (bossimbacon):

thank you !